# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
$\mathcal{E}_{\lambda,s}(\phi v_{j})-\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}$ $=\langle\phi v_{j},\mathcal{P}_{s}(\phi v_{j})\rangle_{2}-\langle v_{j},\mathcal{P}_{s}(\phi^{2}v_{j})\rangle_{2}$
$=\langle v_{j},\phi\mathcal{P}_{s}(\phi v_{j})-P_{s}(\phi^{2}v_{j})\rangle_{2}=-\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}.$
@@
$|\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}|\leq\|v_{j}\|_{H^{s}}\,\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}\leq\|v_{j}\|_{\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})}\,\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}.$
By Lemma 4.7 with $f=\phi v_{j}$,
$\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}\leq C_{\phi}\,\|\phi v_{j}\|_{\tilde{H}^{s-1}}.$
Let $A:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{1/2}$, so that $\|u\|_{\widetilde{H}^{t}}=\|A^{t}u\|_{L^{2}}$ for all $t\in\mathbb{R}$. Let $\{v_{j}\}\subset\widetilde{H}^{s}(\mathbb{H}^{n})$ be bounded and set $f_{j}:=A^{s}v_{j}\in L^{2}(\mathbb{H}^{n})$. Then
$\|\phi v_{j}\|_{\widetilde{H}^{s-1}}=\|A^{s-1}(\phi v_{j})\|_{L^{2}}=\|K_{\phi}f_{j}\|_{L^{2}},$
@@
$K_{\phi}:=A^{\,s-1}M_{\phi}A^{-s}=A^{\,s-1}M_{\phi}A^{-(s-1)}\,A^{-1}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}).$
We claim that $K_{\phi}$ is compact on $L^{2}(\mathbb{H}^{n})$. Indeed, $A^{-1}=(\lambda I-\Delta_{\mathbb{H}^{n}})^{-1/2}$ is an elliptic pseudodifferential operator of order 1-1, hence
$A^{-1}:L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(\mathbb{H}^{n})\quad\text{bounded}.$
@@
$B_{\phi}:=A^{\,s-1}M_{\phi}A^{-(s-1)}\in\Psi^{0}(\mathbb{H}^{n}),\qquad\text{so that}\qquad K_{\phi}=B_{\phi}\,A^{-1}.$
Choose $\chi\in C_{c}^{\infty}(\mathbb{H}^{n})$ such that $\chi\equiv 1$ on a neighbourhood of $\operatorname{supp}\phi$, and set $K:=\operatorname{supp}\chi\Subset\mathbb{H}^{n}$. Since $M_{\phi}=\chi M_{\phi}$, we have
$B_{\phi}=A^{\,s-1}\chi M_{\phi}A^{-(s-1)}=(\chi A^{\,s-1}\chi)\,M_{\phi}\,A^{-(s-1)}.$
@@
$K_{\phi}f=B_{\phi}(A^{-1}f)=(\chi A^{\,s-1}\chi)\,M_{\phi}\,A^{-s}f,$
and the operator $\chi A^{\,s-1}\chi$ is properly supported with Schwartz kernel supported in $K\times K$. Therefore,
$\operatorname{supp}(K_{\phi}f)\subset K,\qquad\forall\,f\in L^{2}(\mathbb{H}^{n}).$
Combining the order 0 boundedness of $B_{\phi}$ with the regularizing property of $A^{-1}$, we obtain a bounded map
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}_{\mathrm{loc}}(\mathbb{H}^{n}).$
Together with $\operatorname{supp}(K_{\phi}f)\subset K$, this yields
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(K)\quad\text{bounded}.$
@@
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow L^{2}(\mathbb{H}^{n})$
is compact. Since $v_{j}\rightharpoonup 0$ weakly in $\tilde{H}^{s}(\mathbb{H}^{n})=\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, $f_{j}\rightharpoonup 0$ weakly in $L^{2}(\mathbb{H}^{n})$, it follows that $K_{\phi}f_{j}\to 0$ strongly in $L^{2}$. Hence
$\|\phi v_{j}\|_{\tilde{H}^{s-1}(\mathbb{H}^{n})}\longrightarrow 0,\qquad j\to\infty.$
Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, it follows that $\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}\to 0$, proving the claim.
@@
Proof of Theorem 1.8. Since $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, we have $H_{n,s,p}(\lambda)>0$. Therefore, it suffices to establish the existence of minimizers for $H_{n,s,p}(\lambda)$, see (4.6).
@@
$I_{\lambda,s}(u_{j})\to H_{n,s}(\lambda).$
@@
where $o(1)\to 0$ as $j\to\infty$ uniformly for $\varphi$ in bounded sets. Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, up to a subsequence,
@@
$\int_{\mathbb{H}^{n}}|\phi v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0\qquad\text{as }j\to\infty.$ (4.22)
@@
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1).$ (4.23)
By Lemma 4.8,
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\phi v_{j})+o(1).$ (4.24)
Combining (4.23)–(4.24) we obtain
$\mathcal{E}_{\lambda,s}(\phi v_{j})=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1).$ (4.25)
@@
$\leq\|\phi v_{j}\|_{p+1}^{2}\left(\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\right)^{\!\frac{p-1}{p+1}}+o(1).$
@@
We now deduce a nontrivial concentration property near $(1,0)\in\mathbb{R}\times\mathbb{R}^{n-1}$. For every $R>0$,
$\liminf_{j\to\infty}\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}>0,$ (4.27)
where $B((1,0),R)$ is the Euclidean ball of radius $R$ centered at $(1,0)$.
@@
$\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0.$ (4.28)
@@
$\int_{B((1,0),1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}=\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}+\int_{A}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0,$
which contradicts the normalization (4.20). Hence (4.28) is impossible, and (4.27) follows. This is impossible if $p+1<2_{s}^{*}$, because we assumed $v=0.$ Next we prove the case of $p+1=2_{s}^{*}.$
Fix $R_{0}\in(0,1)$ and $\psi\in C_{c}^{\infty}(B((1,0),R_{0}))$ with $\psi\equiv 1$ on $B((1,0),R_{0}/2)$ and $0\leq\psi\leq 1$. Similarly, we have
$\liminf_{j\to\infty}\int_{B((1,0),R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\geq H_{n,s}(\lambda)^{\frac{p+1}{p-1}},$ (4.29)
@@
$\int_{\mathbb{H}^{n}\setminus B((1,0),R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\rightarrow 0.$ (4.30)
Proceeding as in (4.23)–(4.24) (with $\phi$ replaced by $\psi$), we obtain
@@
where in the last equality we used $\operatorname{supp}\psi\subset B((1,0),R_{0})$ together.
@@
which contradict with $H_{n,s}(\lambda)<H_{n,s}(0),\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]$. Therefore $v\not\equiv 0$, we complete the proof. $\Box$
@@
$A_{s}:\ C_{c}^{\infty}(\mathbb{B}^{n})\longrightarrow\mathcal{D}^{\prime}(\mathbb{B}^{n})$
@@
$\langle A_{s}f,\varphi\rangle:=\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x),\qquad f,\varphi\in C_{c}^{\infty}(\mathbb{B}^{n}).$
In particular, $A_{s}$ is linear by construction.
@@
AA is continuous: if $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$, then for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$,
$\langle Af_{j},\varphi\rangle\longrightarrow\langle Af,\varphi\rangle.$
Proof: Let $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$. Then there exists a compact set $K\Subset\mathbb{B}^{n}$ such that $\operatorname{supp}f_{j}\subset K$ for all $j$, and for every integer $\ell\geq 0$,
@@
$\|f_{j}-f\|_{H^{m}(\mathbb{B}^{n})}\longrightarrow 0.$
@@
$\bigl|\langle Af_{j}-Af,\varphi\rangle\bigr|=\left|\int_{\mathbb{B}^{n}}\bigl(\widetilde{\mathcal{P}}_{s}(f_{j}-f)\bigr)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)\right|\leq\|\widetilde{\mathcal{P}}_{s}(f_{j}-f)\|_{L^{2}}\,\|\varphi\|_{L^{2}}\longrightarrow 0.$
This proves $\langle Af_{j},\varphi\rangle\to\langle Af,\varphi\rangle$ for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$, i.e. $Af_{j}\rightharpoonup Af$ in $\mathcal{D}^{\prime}(\mathbb{B}^{n})$. ∎
Therefore all assumptions of the Schwartz kernel theorem corresponding to smooth manifolds are satisfied, and there exists a unique distribution $K_{s}\in\mathcal{D}^{\prime}(\mathbb{B}^{n}\times\mathbb{B}^{n})$ such that
@@
Recall that $T_{x}$ is an isometry of $(\mathbb{B}^{n},g_{\mathbb{B}})$ and satisfies $T_{x}(x)=0$, see subsection 2.1. Hence,
@@
$\widehat{\kappa_{s}}(\beta)=\widetilde{m}_{s}(\beta)\qquad\beta\in\mathbb{R}.$ (4.32)
@@
Then $\Phi_{\beta}(0)=1,$ $\Phi_{\beta}(x)$ is ridial and has the explicit Legendre representation
$\Phi_{\beta}(x):=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh d(x,0))^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+i\beta}^{\,\frac{2-n}{2}}(\cosh d(x,0)),$
@@
where $_{2}F_{1}(a,b;c;z)$ is hypergeometric function
$_{2}F_{1}(a,b;c;z)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\,\frac{z^{k}}{k!},\qquad c\neq 0,-1,-2,\dots$
@@
Applying the distributional inversion formula for $\kappa_{s}$ (4.33) yields
@@
$\Phi_{\beta}(r)=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh r)^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+i\beta}^{\,\frac{2-n}{2}}(\cosh r).$
@@
We use the Harish–Chandra expansion from [28, Theorem 3.1] (for rank one) which, in our notation, reads
$\Phi_{\beta}(r)=c(\beta)e^{(i\beta-\rho)r}\phi_{\beta}(r)+c(-\beta)e^{(-i\beta-\rho)r}\phi_{-\beta}(r),\qquad\phi_{\beta}(r)=\sum_{k=0}^{\infty}\Gamma_{k}(\beta)e^{-2kr},$ (4.36)
@@
$J_{+}(r):=\int_{\mathbb{R}}e^{i\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{\beta}(r)}{\overline{c(\beta)}}\,d\beta,\qquad J_{-}(r):=\int_{\mathbb{R}}e^{-i\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{-\beta}(r)}{\overline{c(-\beta)}}\,d\beta.$

### v2

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
$\mathcal{E}_{\lambda,s}(\psi v_{j})-\langle v_{j},\psi^{2}v_{j}\rangle_{\lambda,s}$ $=\langle\psi v_{j},\mathcal{P}_{s}(\psi v_{j})\rangle_{2}-\langle v_{j},\mathcal{P}_{s}(\psi^{2}v_{j})\rangle_{2}$
$=\langle v_{j},\psi\mathcal{P}_{s}(\psi v_{j})-P_{s}(\psi^{2}v_{j})\rangle_{2}=-\langle v_{j},[\mathcal{P}_{s},\psi](\psi v_{j})\rangle_{2}.$
@@
$|\langle v_{j},[\mathcal{P}_{s},\psi](\psi v_{j})\rangle_{2}|\leq\|v_{j}\|_{H^{s}}\,\|[\mathcal{P}_{s},\psi](\psi v_{j})\|_{\tilde{H}^{-s}}\leq\|v_{j}\|_{\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})}\,\|[\mathcal{P}_{s},\psi](\psi v_{j})\|_{\tilde{H}^{-s}}.$
By Lemma 4.7 with $f=\psi v_{j}$,
$\|[\mathcal{P}_{s},\psi](\psi v_{j})\|_{\tilde{H}^{-s}}\leq C_{\psi}\,\|\psi v_{j}\|_{\tilde{H}^{s-1}}.$
Let $\mathcal{A}_{\lambda}:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{1/2}$, so that $\|u\|_{\widetilde{H}^{t}}=\|\mathcal{A}_{\lambda}^{t}u\|_{L^{2}}$ for all $t\in\mathbb{R}$. Let $\{v_{j}\}\subset\widetilde{H}^{s}(\mathbb{H}^{n})$ be bounded and set $f_{j}:=\mathcal{A}_{\lambda}^{s}v_{j}\in L^{2}(\mathbb{H}^{n})$. Then
$\|\psi v_{j}\|_{\widetilde{H}^{s-1}}=\|\mathcal{A}_{\lambda}^{s-1}(\psi v_{j})\|_{L^{2}}=\|{\bf K}_{\psi}f_{j}\|_{L^{2}},$
@@
${\bf K}_{\psi}:=\mathcal{A}_{\lambda}^{\,s-1}M_{\psi}\mathcal{A}_{\lambda}^{-s}=\mathcal{A}_{\lambda}^{\,s-1}M_{\psi}\mathcal{A}_{\lambda}^{-(s-1)}\,\mathcal{A}_{\lambda}^{-1}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}).$
We claim that ${\bf K}_{\psi}$ is compact on $L^{2}(\mathbb{H}^{n})$. Indeed, $\mathcal{A}_{\lambda}^{-1}=(\lambda I-\Delta_{\mathbb{H}^{n}})^{-1/2}$ is an elliptic pseudodifferential operator of order 1-1, hence
$\mathcal{A}_{\lambda}^{-1}:L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(\mathbb{H}^{n})\quad\text{bounded}.$
@@
${\bf B}_{\psi}:=\mathcal{A}_{\lambda}^{\,s-1}M_{\psi}\mathcal{A}_{\lambda}^{-(s-1)}\in\Psi^{0}(\mathbb{H}^{n}),\qquad\text{so that}\qquad{\bf K}_{\psi}={\bf B}_{\psi}\,\mathcal{A}_{\lambda}^{-1}.$
Choose $\chi\in C_{c}^{\infty}(\mathbb{H}^{n})$ such that $\chi\equiv 1$ on a neighbourhood of $\operatorname{supp}\psi$, and set $K:=\operatorname{supp}\chi\Subset\mathbb{H}^{n}$. Since $M_{\psi}=\chi M_{\psi}$, we have
${\bf B}_{\psi}=\mathcal{A}_{\lambda}^{\,s-1}\chi M_{\psi}\mathcal{A}_{\lambda}^{-(s-1)}=(\chi\mathcal{A}_{\lambda}^{\,s-1}\chi)\,M_{\psi}\,\mathcal{A}_{\lambda}^{-(s-1)}.$
@@
${\bf K}_{\psi}f={\bf B}_{\psi}(\mathcal{A}_{\lambda}^{-1}f)=(\chi\mathcal{A}_{\lambda}^{\,s-1}\chi)\,M_{\psi}\,\mathcal{A}_{\lambda}^{-s}f,$
and the operator $\chi\mathcal{A}_{\lambda}^{\,s-1}\chi$ is properly supported with Schwartz kernel supported in $K\times K$. Therefore,
$\operatorname{supp}({\bf K}_{\psi}f)\subset K,\qquad\forall\,f\in L^{2}(\mathbb{H}^{n}).$
Combining the order 0 boundedness of ${\bf B}_{\psi}$ with the regularizing property of $\mathcal{A}_{\lambda}^{-1}$, we obtain a bounded map
${\bf K}_{\psi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}_{\mathrm{loc}}(\mathbb{H}^{n}).$
Together with $\operatorname{supp}({\bf K}_{\psi}f)\subset K$, this yields
${\bf K}_{\psi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(K)\quad\text{bounded}.$
@@
${\bf K}_{\psi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow L^{2}(\mathbb{H}^{n})$
is compact. Since $v_{j}\rightharpoonup 0$ weakly in $\tilde{H}^{s}(\mathbb{H}^{n})=\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, $f_{j}\rightharpoonup 0$ weakly in $L^{2}(\mathbb{H}^{n})$ as $j\to+\infty$, it follows that ${\bf K}_{\psi}f_{j}\to 0$ strongly in $L^{2}$. Hence
$\|\psi v_{j}\|_{\tilde{H}^{s-1}(\mathbb{H}^{n})}\longrightarrow 0\quad{\rm as}\ \,j\to+\infty.$
Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, it follows that $\langle v_{j},[\mathcal{P}_{s},\psi](\psi v_{j})\rangle_{2}\to 0$, proving the claim.
@@
In this subsection, we complete the proofs of Theorems 1.7 and 1.8.
Proof of Theorem 1.8. Since $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, we have $H_{n,s,p}(\lambda)>0$ by (4.6). Therefore, it suffices to establish the existence of minimizers for $H_{n,s,p}(\lambda)$.
@@
$I_{\lambda,s}(u_{j})\to H_{n,s}(\lambda)\quad{\rm as}\ j\to+\infty.$
@@
where $o(1)\to 0$ as $j\to\infty$ uniformly for $\varphi$ in bounded sets. Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$ and as $j\to\infty$, up to a subsequence,
@@
$\int_{\mathbb{H}^{n}}|\phi v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0\quad\text{as }j\to\infty.$ (4.22)
@@
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1)\quad{\rm as}\ j\to+\infty.$ (4.23)
By Lemma 4.8, one has
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\phi v_{j})+o(1),$ (4.24)
which, combining (4.23)–(4.24), implies that
$\mathcal{E}_{\lambda,s}(\phi v_{j})=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1)\quad{\rm as}\ j\to+\infty.$ (4.25)
@@
$\leq\|\phi v_{j}\|_{p+1}^{2}\left(\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\right)^{\!\frac{p-1}{p+1}}+o(1)\qquad{\rm as}\ j\to+\infty.$
@@
We now deduce a nontrivial concentration property near $e_{1}:=(1,0,\cdots,0)\in\mathbb{R}\times\mathbb{R}^{n-1}$. For every $R>0$,
$\liminf_{j\to\infty}\int_{B(e_{1},R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}>0,$ (4.27)
where $B(e_{1},R)$ is the Euclidean ball of radius $R$ centered at $(1,0)$.
@@
$\int_{B(e_{1},R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0\quad{\rm as}\ \,j\to+\infty.$ (4.28)
@@
$\int_{B(e_{1},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}=\int_{B(e_{1},R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}+\int_{A}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0,$
which contradicts the normalization (4.20). Hence (4.28) fails and (4.27) follows. Then it is impossible for $p+1<2_{s}^{*}$, because we assumed $v=0.$
Next, we prove the case of $p+1=2_{s}^{*}.$
Fix $R_{0}\in(0,1)$ and $\psi\in C_{c}^{\infty}(B(e_{1},R_{0}))$ with $\psi\equiv 1$ on $B(e_{1},R_{0}/2)$ and $0\leq\psi\leq 1$. Similarly, we have
$\liminf_{j\to\infty}\int_{B(e_{1},R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\geq H_{n,s}(\lambda)^{\frac{p+1}{p-1}},$ (4.29)
@@
$\int_{\mathbb{H}^{n}\setminus B(e_{1},R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\rightarrow 0.$ (4.30)
Proceeding as in (4.23)–(4.24) (with $\phi$ replaced by $\psi$), we obtain that
@@
where in the last equality we used $\operatorname{supp}\psi\subset B(e_{1},R_{0})$ together.
@@
which contradict with $H_{n,s}(\lambda)<H_{n,s}(0),\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]$. Therefore $v\not\equiv 0$, which ends the proof. $\Box$
@@
In this subsection, we analyze the monotonicity, attainability, and strict-gap regimes of the Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, and prove Theorem 1.5 and Theorem 1.6. The key step is to show the threshold phenomenon that, once $\lambda$ is above the bottom of the corresponding spectrum, the level drops instantly to -$\infty$. For nonlocal operators, this requires delicate control of far-interaction terms. To obtain this off-diagonal decay, we combine the Schwartz kernel theorem (yielding a distributional kernel representation) with the Harish–Chandra asymptotic expansion.
@@
${\bf A}_{s}:\ C_{c}^{\infty}(\mathbb{B}^{n})\longrightarrow\mathcal{D}^{\prime}(\mathbb{B}^{n})$
@@
$\langle{\bf A}_{s}f,\varphi\rangle:=\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)\quad{\rm for}\ \,f,\varphi\in C_{c}^{\infty}(\mathbb{B}^{n}).$
In particular, ${\bf A}_{s}$ is linear by construction.
@@
The operator ${\bf A}_{s}$ is continuous in the following sense: if $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$, then for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$,
$\langle{\bf A}_{s}f_{j},\varphi\rangle\longrightarrow\langle{\bf A}_{s}f,\varphi\rangle\quad{\rm as}\ j\to+\infty.$
Proof: Let $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$ as $j\to+\infty$, then there exists a compact set $K\Subset\mathbb{B}^{n}$ such that $\operatorname{supp}f_{j}\subset K$ for all $j$, and for every integer $\ell\geq 0$,
@@
$\|f_{j}-f\|_{H^{m}(\mathbb{B}^{n})}\longrightarrow 0\quad{\rm as}\ j\to+\infty.$
@@
$\bigl|\langle{\bf A}_{s}f_{j}-{\bf A}_{s}f,\varphi\rangle\bigr|=\left|\int_{\mathbb{B}^{n}}\bigl(\widetilde{\mathcal{P}}_{s}(f_{j}-f)\bigr)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)\right|\leq\|\widetilde{\mathcal{P}}_{s}(f_{j}-f)\|_{L^{2}}\,\|\varphi\|_{L^{2}}\longrightarrow 0.$
This proves $\langle{\bf A}_{s}f_{j},\varphi\rangle\to\langle{\bf A}_{s}f,\varphi\rangle$ for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$, i.e. ${\bf A}_{s}f_{j}\rightharpoonup{\bf A}_{s}f$ in $\mathcal{D}^{\prime}(\mathbb{B}^{n})$. ∎
Now all assumptions of the Schwartz kernel theorem corresponding to smooth manifolds are satisfied, and there exists a unique distribution $K_{s}\in\mathcal{D}^{\prime}(\mathbb{B}^{n}\times\mathbb{B}^{n})$ such that
@@
Recall that the Möbius transformation $T_{x}$ is an isometry of $(\mathbb{B}^{n},g_{\mathbb{B}})$ and satisfies $T_{x}(x)=0$. Hence,
@@
$\widehat{\kappa_{s}}(\beta)=\widetilde{m}_{s}(\beta)\quad{\rm for}\ \,\beta\in\mathbb{R}.$ (4.32)
@@
where $h_{-\beta,\theta}$ is defined in (2.2). Then $\Phi_{\beta}(0)=1,$ $\Phi_{\beta}(x)$ is ridial and has the explicit Legendre representation
$\Phi_{\beta}(x):=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh d(x,0))^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+{\bf i}\beta}^{\,\frac{2-n}{2}}(\cosh d(x,0)),$
@@
where $_{2}F_{1}(a,b;c;z)$ is a hypergeometric function
$_{2}F_{1}(a,b;c;z)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\,\frac{z^{k}}{k!},\qquad c\neq 0,-1,-2,\dots.$
The next proposition provides the crucial off-diagonal exponential decay estimate for $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, which controls long-range interactions between separated supports.
@@
Applying the distributional inversion formula (4.33) for $\kappa_{s}$ yields
@@
$\Phi_{\beta}(r)=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh r)^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+{\bf i}\beta}^{\,\frac{2-n}{2}}(\cosh r).$
@@
We use the Harish–Chandra expansion from [40, Theorem 3.1] (for rank one) which, in our notation, reads
$\Phi_{\beta}(r)=c(\beta)e^{({\bf i}\beta-\rho)r}\phi_{\beta}(r)+c(-\beta)e^{(-{\bf i}\beta-\rho)r}\phi_{-\beta}(r),\qquad\phi_{\beta}(r)=\sum_{k=0}^{\infty}\Gamma_{k}(\beta)e^{-2kr},$ (4.36)
@@
$J_{+}(r):=\int_{\mathbb{R}}e^{{\bf i}\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{\beta}(r)}{\overline{c(\beta)}}\,d\beta,\qquad J_{-}(r):=\int_{\mathbb{R}}e^{-{\bf i}\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{-\beta}(r)}{\overline{c(-\beta)}}\,d\beta.$